### v1

@@ 3. $(\mathcal{S}_{3})$
Let $F\in M(\Delta_{3}(W_{n}))$ such that $F\in\mathcal{T}_{s}$ and ${F}^{c}=\{\alpha_{s}\}\sqcup\{s_{1},s_{2}\}$. Suppose there exists $\mu\in F$ such that $\alpha_{s}<_{\mathcal{O}}\mu$ and $s_{1}<\mu$. Let ${F^{\prime}}^{c}=({F}^{c}\setminus\{s_{2}\})\sqcup\{\mu\}$. If either $F^{\prime}\notin M(\Delta_{3}(W_{n}))$, or $F^{\prime}\in\mathcal{D}$ whenever $F^{\prime}\in M(\Delta_{3}(W_{n}))$, then $F$ is not a spanning facet.
@@
Assume that either $F^{\prime}\notin M(\Delta_{3}(W_{n}))$, or $F^{\prime}\in\mathcal{D}$ whenever $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Suppose $F$ is a spanning facet. We have $\mu\in F$. By Remark 3.10, there exists $\tilde{F}\in M(\Delta_{3}(W_{n}))$ such that $\tilde{F}\prec F$ and $\tilde{F}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{\mu\}$ for some $\nu\in F^{c}$. Moreover, $F\notin\mathcal{D}$ by Remark 3.15, and $s_{2}=n-1$ by Proposition 3.17.
If $\nu=\alpha_{s}$, then $\alpha_{s}<_{\mathcal{O}}\mu$ implies that $F\prec\tilde{F}$ by Proposition 3.16, a contradiction. Also, if $\nu=s_{1}$, then $\alpha_{s}<_{\mathcal{O}}\mu$ implies that $\tilde{F}^{c}=\{\alpha_{s}\}\sqcup\{\mu,n-1\}$. It follows that $\tilde{F}\in\mathcal{T}_{s}$, and since $s_{1}<\mu$, $F\ll\tilde{F}$. Hence $F\notin\mathcal{D}$ implies that $F\prec\tilde{F}$ by Definition 3.4 (i) and (iii), again a contradiction. So, $\nu=s_{2}=n-1$. Then $\tilde{F}=F^{\prime}$.
If $F^{\prime}\notin M(\Delta_{3}(W_{n}))$, then this implies that $\tilde{F}$ is not a facet, which is a contradiction.
If $F^{\prime}\in M(\Delta_{3}(W_{n}))$, then by our assumption, $F^{\prime}\in\mathcal{D}$. Now, since $\alpha_{s}<_{\mathcal{O}}\mu$ and $s_{1}<\mu$, we get $F^{\prime c}=\{\alpha_{s}\}\sqcup\{s_{1},\mu\}$ and thus $F^{\prime}\in\mathcal{T}_{s}$. Further, since $F\notin\mathcal{D}$, $F\prec F^{\prime}$ by Definition 3.4 (iii), which is a contradiction.
@@
Let $F\in M(\Delta_{3}(W_{n}))$ be a spanning facet. Then $F\in\mathcal{S}$.
@@
    We show that $F$ satisfies $(\mathcal{S}_{1})$. For this, we need to show that $p_{1}\in V(W_{n})\setminus(\{0,1,2,3,4,\ldots,2m-4\}\sqcup\{n-1\})$. We have $\alpha_{p}=3<m$ (as $m\geq 5$) and $\alpha_{p}<_{\mathcal{O}}p_{1}$. By Remark 3.1 (i), either $p_{1}<3$ or $p_{1}\geq 2m-3$. Thus $p_{1}\notin\{3,4,\ldots,2m-4\}$.
    Suppose $p_{1}<3$. Then $W_{n}[F^{c}]$ is disconnected implies that $p_{1}\in\{0,2\}$. It follows that either $F^{c}=\{n-1,(n-1)+1\ (\text{mod \ $n$}),(n-1)+4\ (\text{mod \ $n$})\}$ or $F^{c}=\{n-1,(n-1)+3\ (\text{mod \ $n$}),(n-1)+4\ (\text{mod \ $n$})\}$, which is a contradiction to Proposition 3.14. Hence, $p_{1}\geq 2m-3$. Since $p_{1}<p_{2}=n-1$, $p_{1}\neq n-1$. This means that $p_{1}\in V(W_{n})\setminus(\{0,1,2,3,4,\ldots,2m-4\}\sqcup\{n-1\})$.
@@
    In this case, we show that $p_{1}\in V(W_{n})\setminus(\{\alpha_{p}-4,\alpha_{p}-3,\alpha_{p}-2\}\sqcup\{\alpha_{p},\alpha_{p}+1,\ldots,2m-\alpha_{p}-1\}\sqcup\{n-1\})$, which implies that $F$ satisfies $(\mathcal{S}_{2})$.
@@
    Suppose $p_{1}\in\{\alpha_{p}-4,\alpha_{p}-3\}$. We have $p_{1}+2\leq\alpha_{p}-1\leq m-3<n-1=p_{2}$. Thus $p_{1}+2\in F$. Clearly, $p_{1}<p_{1}+2$. Moreover, observe that $\alpha_{p}<_{\mathcal{O}}p_{1}+2$, and $W_{n}[(F^{c}\setminus\{p_{2}\})\sqcup\{p_{1}+2\}]=W_{n}[\{p_{1},p_{1}+2,\alpha_{p}\}]$ is connected. Thus $F^{\prime}\notin M(\Delta_{3}(W_{n}))$, where $F^{\prime c}=(F^{c}\setminus\{p_{2}\})\sqcup\{p_{1}+2\}$. It follows from Proposition 3.18 that $F$ is not a spanning facet, a contradiction. Therefore $p_{1}\notin\{\alpha_{p}-4,\alpha_{p}-3\}$.
    Now, suppose $p_{1}=\alpha_{p}-2$. Since $W_{n}[F^{c}]$ is disconnected, $p+1=\alpha_{p}-1\in F$. We have $\alpha_{p}<_{\mathcal{O}}\alpha_{p}-1=p+1$ and $p_{1}<p_{1}+1$. Moreover, $W_{n}[(F^{c}\setminus\{p_{2}\})\sqcup\{p_{1}+1\}]=W_{n}[\{\alpha_{p}-2,\alpha_{p}-1,\alpha_{p}\}]$ is connected. Hence, using Proposition 3.18, $F$ is not a spanning facet, a contradiction. Thus, $p_{1}\neq\alpha_{p}-2$.
    Therefore, $p_{1}\in V(W_{n})\setminus(\{\alpha_{p}-4,\alpha_{p}-3,\alpha_{p}-2,\alpha_{p},\alpha_{p}+1,\ldots,2m-\alpha_{p}-1\}\sqcup\{n-1\})$.
@@
    We show that $p_{1}\in V(W_{n})\setminus(\{\alpha_{p}-4,\alpha_{p}-3,\alpha_{p}-2,\alpha_{p}-1,\alpha_{p},\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\ (\text{mod \ $n$})\}\sqcup\{y\})$, which implies that $F$ satisfies $(\mathcal{S}_{3})$ (as $\min\{2m-\alpha_{p},\alpha_{p}-4\}=\alpha_{p}-4$).
@@
    Suppose $p_{1}=y$. We have $\alpha_{p}\leq m\leq n-4$ (as $n\geq 9$). So $y\neq 1$. Since $p_{1}\neq n-1,$ we get $y=0$. Clearly, if $\alpha_{p}\neq n-4$, then $y\neq 0$. Thus $\alpha_{p}=n-4$. This means that $n=9$ and $\alpha_{p}=m$, thereby implying that $F^{c}=\{m,m+3,m+4\ (\text{mod $n$)}\}$. Then $F$ is not a spanning facet by Proposition 3.14. This is a contradiction. So $p_{1}\neq y$.
    We now show that if $p_{1}$$\in\{\alpha_{p}-4,\alpha_{p}-3,\alpha_{p}-2,\alpha_{p}-1,\alpha_{p}+1,\alpha_{p}+2\}$, then there exists $\mu\in F$ such that $\alpha_{p}<_{\mathcal{O}}\mu$, $p_{1}<\mu$ and $W_{n}[(F^{c}\setminus\{p_{2}\})\sqcup\{\mu\}]$ is connected. Then Proposition 3.18 implies that $F$ is not a spanning facet, which is a contradiction.
@@
      We have $p_{1}+2\leq\alpha_{p}-1\leq m-1<n-1=p_{2}$. It follows that $p_{1}+2\in F$. Take, $\mu=p_{1}+2$. Then $\alpha_{p}<_{\mathcal{O}}\mu$, $p_{1}<\mu$ and $W_{n}[(F^{c}\setminus\{p_{2}\})\sqcup\{\mu\}]=W_{n}[\{p_{1},p_{1}+2,\alpha_{p}\}]$ is connected.
@@
      Since $p_{1}+1=\alpha_{p}-1\leq m-1<n-1=p_{2}$, we have $p_{1}+1\in F$. Take $\mu=p_{1}+1$. Then $\alpha_{p}<_{\mathcal{O}}\mu$, $p_{1}<\mu$ and $W_{n}[(F^{c}\setminus\{p_{2}\})\sqcup\{\mu\}]=W_{n}[\{\alpha_{p}-2,\alpha_{p}-1,\alpha_{p}\}]$ is connected.
@@
      Here, $p_{1}+3=\alpha_{p}+2\leq m+2<n-1=p_{2}$, which implies that $p_{1}+3\in F$. Take $\mu=p_{1}+3$. Then $\alpha_{p}<_{\mathcal{O}}\mu$, $p_{1}<\mu$ and $W_{n}[(F^{c}\setminus\{p_{2}\})\sqcup\{\mu\}]=W_{n}[\{\alpha_{p}-1,\alpha_{p},\alpha_{p}+2\}]$ is connected.
@@
      In this case, $\alpha_{p}<_{\mathcal{O}}p_{1}$ implies that $\alpha_{p}=m$. We have $p_{1}+1=m+2<n-1$. So $p_{1}+1\in F$. Take, $\mu=p_{1}+1$. We have $\alpha_{p}=m<_{\mathcal{O}}m+2=p_{1}+1=\mu$ and $p_{1}<\mu$. Clearly, $W_{n}[(F^{c}\setminus\{p_{2}\})\sqcup\{\mu\}]=W_{n}[\{\alpha_{p},\alpha_{p}+1,\alpha_{p}+2\}]$ is connected.
@@
      Here, since $W_{n}[F^{c}]$ is disconnected, we have $n\geq 10$. It follows that $p_{1}+1=\alpha_{p}+3\leq m+3<n-1$, and thus $p_{1}+1\in F$. Take $\mu=p_{1}+1$. Observe that $\alpha_{p}<_{\mathcal{O}}\mu$ and $W_{n}[(F^{c}\setminus\{p_{2}\})\sqcup\{\mu\}]=W_{n}[\{\alpha_{p},\alpha_{p}+2,\alpha_{p}+3\}]$ is connected.
@@
    Now, suppose $p_{1}=\alpha_{p}+3\ (\text{mod \ $n$})$. Since $\alpha_{p}\leq m$, we have $\alpha_{p}+3\ (\text{mod \ $n$})=\alpha_{p}+3$. We first assume that $n=9$. Then, $m+3=n-1$. Therefore, if $\alpha_{p}=m-1$, then $F^{c}=\{\alpha_{p},\alpha_{p}+3,\alpha_{p}+4\}$, which implies that $F$ is not a spanning facet by Proposition 3.14. This is a contradiction. So $\alpha_{p}=m$. Then $p_{1}=m+3=n-1=p_{2}$, which is again a contradiction.
    Thus, we assume that $n\geq 10$. It follows that $p_{1}+1=\alpha_{p}+4\leq m+4\leq n-1$. If $p_{1}+1=n-1$, then $\alpha_{p}=m$ and $F^{c}=\{m,m+3,m+4\}$. By Proposition 3.14, it follows that $F$ is not a spanning facet, a contradiction. Hence $p_{1}+1<n-1$ and thus, $p_{1}+1\in F$. We have $\alpha_{p}<_{\mathcal{O}}p_{1}+1$ and $p_{1}<p_{1}+1$. Let $F^{\prime c}=(F^{c}\setminus\{p_{2}\})\sqcup\{p_{1}+1\}$. If $F^{\prime}\notin M(\Delta_{3}(W_{n}))$, then Proposition 3.18 implies that $F$ is not a spanning facet, which is a contradiction. So $F^{\prime}\in M(\Delta_{3}(W_{n}))$. Then $F^{\prime c}=\{\alpha_{p}\}\sqcup\{\alpha_{p}+3,\alpha_{p}+4\}$ satisfies $(\mathcal{D}_{4})$, and thus $F^{\prime}\in\mathcal{D}$. Hence, Proposition 3.18 implies that $F$ is not a spanning facet, again a contradiction. So $p_{1}\neq\alpha_{p}+3\text{ (mod \ $n$})$.
    Therefore, $p_{1}\in V(W_{n})\setminus(\{\alpha_{p}-4,\alpha_{p}-3,\alpha_{p}-2,\alpha_{p}-1,\alpha_{p},\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\ (\text{mod \ $n$})\}\sqcup\{y\})$.
@@
    We show that $F$ satisfies $(\mathcal{S}_{3})$. For this, we show that $p_{1}\in V(W_{n})\setminus(\{m-3,m-2,m-1,m,m+1,m+2,m+3,m+4\ (\text{mod $n$})\}\sqcup\{y\})$ (as $\min\{2m-\alpha_{p},\alpha_{p}-4\}=\alpha_{p}-4=m-3$). We have $F^{c}=\{m+1\}\sqcup\{p_{1},p_{2}\}$, where $p_{2}=n-1$.
    Suppose $p_{1}=y$. Since $p_{1}\neq p_{2}=n-1$, we get $y\in\{0,1\}$. If $y=0$, then $m+1=n-4$. Hence $p_{2}=n-1=m+4$, which implies that $F^{c}=\{m+1,(m+1)+3\text{ (mod \ $n$}),(m+1)+4\text{ (mod \ $n$})\}$. Thus, $F$ is not a spanning facet by Proposition 3.14. This is a contradiction. So $y=1$, and then $m+1>n-4$. This implies that $W_{n}[F^{c}]$ is connected, again a contradiction. Hence $p_{1}\neq y$.
@@
      Here, $p_{1}+1=m-2\in F$. We have $\alpha_{p}=m+1<_{\mathcal{O}}m-2=p_{1}+1$ and $p_{1}<p_{1}+1$. Observe that if $F^{\prime c}=\{m+1\}\sqcup\{m-3,m-2\}$, then $F^{\prime}\in M(\Delta_{3}(W_{n}))$ such that $F^{\prime}$ satisfies $(\mathcal{D}_{3})$ and thus, $F^{\prime}\in\mathcal{D}$. Therefore, Proposition 3.18 implies that $F$ is not a spanning facet, a contradiction. Hence $p_{1}\neq m-3$.
@@
      We have $p_{1}+4=m+2\in F$. Also, $\alpha_{p}=m+1<_{\mathcal{O}}m+2=p_{1}+4$ and $p_{1}<p_{1}+4$. If $F^{\prime c}=\{m+1\}\sqcup\{m-2,m+2\}$, then $F^{\prime}\in M(\Delta_{3}(W_{n}))$ such that $F^{\prime}$ satisfies $(\mathcal{D}_{1})$. Thus $F^{\prime}\in\mathcal{D}$. Therefore, by Proposition 3.18, $F$ is not a spanning facet. This is a contradiction. Hence $p_{1}\neq m-2$.
@@
      Observe that if $n=9$, then $n-1=m+3$; if $n\in\{10,11\}$, then $n-1=m+4$; and if $n\in\{12,13\}$, then $n-1=m+5$. We have $F^{c}=\{m+1\}\sqcup\{p_{1},n-1\}$. Since $W_{n}[F^{c}]$ is disconnected and $p_{1}<n-1$, it follows that if $p_{1}=m+2$, then $n\geq 12$; and if $p_{1}=m+3$, then $n\geq 14$. Therefore, in each case, we have $p_{1}+2<n-1$. This implies that $p_{1}+2\in F$. We have $\alpha_{p}<_{\mathcal{O}}p_{1}+2$, $p_{1}<p_{1}+2$ and $W_{n}[(F^{c}\setminus\{n-1\})\sqcup\{p_{1}+2\}]=W_{n}[\{\alpha_{p},p_{1},p_{1}+2\}]$ is connected. Using Proposition 3.18, we get a contradiction to the fact that $F$ is a spanning facet. Hence $p_{1}\notin\{m+2,m+3\}$.

### v2

@@ 3. $(\mathcal{S}_{3})$
Let $F\in M(\Delta_{3}(C_{n}^{2}))$ such that $F\in\mathcal{T}_{s}$ and ${F}^{c}=\{\alpha_{s}\}\sqcup\{s_{1},s_{2}\}$. Suppose there exists $\mu\in F$ such that $\alpha_{s}<_{\mathcal{O}}\mu$ and $s_{1}<\mu$. Let ${F^{\prime}}^{c}=({F}^{c}\setminus\{s_{2}\})\sqcup\{\mu\}$. If either $F^{\prime}\notin M(\Delta_{3}(C_{n}^{2}))$, or $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$ such that $F^{\prime}\in\mathcal{D}$, then $F$ is not a spanning facet.
@@
Assume that either $F^{\prime}\notin M(\Delta_{3}(C_{n}^{2}))$, or $F^{\prime}\in\mathcal{D}$ whenever $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Suppose $F$ is a spanning facet. We have $\mu\in F$. By Remark 3.11, there exists $\tilde{F}\in M(\Delta_{3}(C_{n}^{2}))$ such that $\tilde{F}\prec F$ and $\tilde{F}^{c}=(F^{c}\setminus\{\nu\})\sqcup\{\mu\}$ for some $\nu\in F^{c}$. Moreover, $F\notin\mathcal{D}$ by Remark 3.16, and $s_{2}=n-1$ by Proposition 3.18.
If $\nu=\alpha_{s}$, then $\alpha_{s}<_{\mathcal{O}}\mu$ implies that $F\prec\tilde{F}$ by Proposition 3.17, a contradiction. Now, if $\nu=s_{1}$, then $\alpha_{s}<_{\mathcal{O}}\mu$ implies that $\tilde{F}^{c}=\{\alpha_{s}\}\sqcup\{\mu,n-1\}$. It follows that $\tilde{F}\in\mathcal{T}_{s}$, and since $s_{1}<\mu$, $F\ll\tilde{F}$. Hence $F\notin\mathcal{D}$ implies that $F\prec\tilde{F}$ by Definition 3.5 (i) and (iii), again a contradiction. So, $\nu=s_{2}=n-1$. Then $\tilde{F}=F^{\prime}$.
If $F^{\prime}\notin M(\Delta_{3}(C_{n}^{2}))$, then this implies that $\tilde{F}\notin M(\Delta_{3}(C_{n}^{2}))$, which is a contradiction.
If $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$, then by our assumption, $F^{\prime}\in\mathcal{D}$. Now, since $\alpha_{s}<_{\mathcal{O}}\mu$ and $s_{1}<\mu$, we get $F^{\prime c}=\{\alpha_{s}\}\sqcup\{s_{1},\mu\}$ and thus $F^{\prime}\in\mathcal{T}_{s}$. Further, since $F\notin\mathcal{D}$, $F\prec F^{\prime}$ by Definition 3.5 (iii), which is a contradiction.
@@
Let $F\in M(\Delta_{3}(C_{n}^{2}))$ be a spanning facet. Then $F\in\mathcal{S}$.
@@
    We show that $F$ satisfies $(\mathcal{S}_{1})$. For this, we need to show that $p_{1}\in V(C_{n}^{2})\setminus(\{0,1,2,3,4,\ldots,2m-4\}\sqcup\{n-1\})$. We have $\alpha_{p}=3<m$ (as $m\geq 5$) and $\alpha_{p}<_{\mathcal{O}}p_{1}$. By Remark 3.1 (i), either $p_{1}<3$ or $p_{1}\geq 2m-3$. Thus $p_{1}\notin\{3,4,\ldots,2m-4\}$.
    Suppose $p_{1}<3$. Then $C_{n}^{2}[F^{c}]$ is disconnected implies that $p_{1}\in\{0,2\}$. It follows that either $F^{c}=\{n-1,(n-1)+1\ (\text{mod \ $n$}),(n-1)+4\ (\text{mod \ $n$})\}$ or $F^{c}=\{n-1,(n-1)+3\ (\text{mod \ $n$}),(n-1)+4\ (\text{mod \ $n$})\}$, which is a contradiction to Proposition 3.15. Hence, $p_{1}\geq 2m-3$. Since $p_{1}<p_{2}=n-1$, $p_{1}\neq n-1$. This means that $p_{1}\in V(C_{n}^{2})\setminus(\{0,1,2,3,4,\ldots,2m-4\}\sqcup\{n-1\})$.
@@
    In this case, we show that $p_{1}\in V(C_{n}^{2})\setminus(\{\alpha_{p}-4,\alpha_{p}-3,\alpha_{p}-2\}\sqcup\{\alpha_{p},\alpha_{p}+1,\ldots,2m-\alpha_{p}-1\}\sqcup\{n-1\})$, which implies that $F$ satisfies $(\mathcal{S}_{2})$.
@@
    Suppose $p_{1}\in\{\alpha_{p}-4,\alpha_{p}-3\}$. We have $p_{1}+2\leq\alpha_{p}-1\leq m-3<n-1=p_{2}$. Thus $p_{1}+2\in F$. Clearly, $p_{1}<p_{1}+2$. Moreover, observe that $\alpha_{p}<_{\mathcal{O}}p_{1}+2$, and $C_{n}^{2}[(F^{c}\setminus\{p_{2}\})\sqcup\{p_{1}+2\}]=C_{n}^{2}[\{p_{1},p_{1}+2,\alpha_{p}\}]$ is connected. Thus $F^{\prime}\notin M(\Delta_{3}(C_{n}^{2}))$, where $F^{\prime c}=(F^{c}\setminus\{p_{2}\})\sqcup\{p_{1}+2\}$. It follows from Proposition 3.19 that $F$ is not a spanning facet, a contradiction. Therefore $p_{1}\notin\{\alpha_{p}-4,\alpha_{p}-3\}$.
    Now, suppose $p_{1}=\alpha_{p}-2$. Since $C_{n}^{2}[F^{c}]$ is disconnected, $p+1=\alpha_{p}-1\in F$. We have $\alpha_{p}<_{\mathcal{O}}\alpha_{p}-1=p+1$ and $p_{1}<p_{1}+1$. Moreover, $C_{n}^{2}[(F^{c}\setminus\{p_{2}\})\sqcup\{p_{1}+1\}]=C_{n}^{2}[\{\alpha_{p}-2,\alpha_{p}-1,\alpha_{p}\}]$ is connected. Hence, using Proposition 3.19, $F$ is not a spanning facet, a contradiction. Thus, $p_{1}\neq\alpha_{p}-2$.
    Therefore, $p_{1}\in V(C_{n}^{2})\setminus(\{\alpha_{p}-4,\alpha_{p}-3,\alpha_{p}-2,\alpha_{p},\alpha_{p}+1,\ldots,2m-\alpha_{p}-1\}\sqcup\{n-1\})$.
@@
    $y:=\begin{cases}n-1&\text{ if }\alpha_{p}<n-4,\\ 0&\text{ if }\alpha_{p}=n-4,\\ 1&\text{ if }\alpha_{p}>n-4.\end{cases}$
    We show that $p_{1}\in V(C_{n}^{2})\setminus(\{\alpha_{p}-4,\alpha_{p}-3,\alpha_{p}-2,\alpha_{p}-1,\alpha_{p},\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\ (\text{mod \ $n$})\}\sqcup\{y\})$, which implies that $F$ satisfies $(\mathcal{S}_{3})$ (as $\min\{2m-\alpha_{p},\alpha_{p}-4\}=\alpha_{p}-4$).
@@
    Suppose $p_{1}=y$. We have $\alpha_{p}\leq m\leq n-4$. So $y\neq 1$. Since $p_{1}\neq n-1,$ we get $y=0$. Clearly, if $\alpha_{p}\neq n-4$, then $y\neq 0$. Thus $\alpha_{p}=n-4$. This means that $n=9$ and $\alpha_{p}=m$, thereby implying that $F^{c}=\{m,m+3,m+4\ (\text{mod $n$)}\}$. Then $F$ is not a spanning facet by Proposition 3.15. This is a contradiction. So $p_{1}\neq y$.
    We now show that if $p_{1}$$\in\{\alpha_{p}-4,\alpha_{p}-3,\alpha_{p}-2,\alpha_{p}-1,\alpha_{p}+1,\alpha_{p}+2\}$, then there exists $\mu\in F$ such that $\alpha_{p}<_{\mathcal{O}}\mu$, $p_{1}<\mu$ and $C_{n}^{2}[(F^{c}\setminus\{p_{2}\})\sqcup\{\mu\}]$ is connected. Then Proposition 3.19 implies that $F$ is not a spanning facet, which is a contradiction.
@@
      We have $p_{1}+2\leq\alpha_{p}-1\leq m-1<n-1=p_{2}$. It follows that $p_{1}+2\in F$. Take, $\mu=p_{1}+2$. Then $\alpha_{p}<_{\mathcal{O}}\mu$, $p_{1}<\mu$ and $C_{n}^{2}[(F^{c}\setminus\{p_{2}\})\sqcup\{\mu\}]=C_{n}^{2}[\{p_{1},p_{1}+2,\alpha_{p}\}]$ is connected.
@@
      Since $p_{1}+1=\alpha_{p}-1\leq m-1<n-1=p_{2}$, we have $p_{1}+1\in F$. Take $\mu=p_{1}+1$. Then $\alpha_{p}<_{\mathcal{O}}\mu$, $p_{1}<\mu$ and $C_{n}^{2}[(F^{c}\setminus\{p_{2}\})\sqcup\{\mu\}]=C_{n}^{2}[\{\alpha_{p}-2,\alpha_{p}-1,\alpha_{p}\}]$ is connected.
@@
      Here, $p_{1}+3=\alpha_{p}+2\leq m+2<n-1=p_{2}$, which implies that $p_{1}+3\in F$. Take $\mu=p_{1}+3$. Then $\alpha_{p}<_{\mathcal{O}}\mu$, $p_{1}<\mu$ and $C_{n}^{2}[(F^{c}\setminus\{p_{2}\})\sqcup\{\mu\}]=C_{n}^{2}[\{\alpha_{p}-1,\alpha_{p},\alpha_{p}+2\}]$ is connected.
@@
      In this case, $\alpha_{p}<_{\mathcal{O}}p_{1}$ implies that $\alpha_{p}=m$. Then $p_{1}+1=m+2<n-1$. So $p_{1}+1\in F$. Take, $\mu=p_{1}+1$. We have $\alpha_{p}=m<_{\mathcal{O}}m+2=p_{1}+1=\mu$ and $p_{1}<\mu$. Clearly, $C_{n}^{2}[(F^{c}\setminus\{p_{2}\})\sqcup\{\mu\}]=C_{n}^{2}[\{\alpha_{p},\alpha_{p}+1,\alpha_{p}+2\}]$ is connected.
@@
      Here, since $C_{n}^{2}[F^{c}]$ is disconnected, we have $n\geq 10$. It follows that $p_{1}+1=\alpha_{p}+3\leq m+3<n-1$, and thus $p_{1}+1\in F$. Take $\mu=p_{1}+1$. Observe that $\alpha_{p}<_{\mathcal{O}}\mu$ and $C_{n}^{2}[(F^{c}\setminus\{p_{2}\})\sqcup\{\mu\}]=C_{n}^{2}[\{\alpha_{p},\alpha_{p}+2,\alpha_{p}+3\}]$ is connected.
@@
    Now, suppose $p_{1}=\alpha_{p}+3\ (\text{mod \ $n$})$. Since $\alpha_{p}\leq m$, we have $\alpha_{p}+3\ (\text{mod \ $n$})=\alpha_{p}+3$. Suppose $n=9$. Then $m+3=n-1$. Therefore, if $\alpha_{p}=m-1$, then $F^{c}=\{\alpha_{p},\alpha_{p}+3,\alpha_{p}+4\}$, which implies that $F$ is not a spanning facet by Proposition 3.15. This is a contradiction. So $\alpha_{p}=m$. Then $p_{1}=m+3=n-1=p_{2}$, which is again a contradiction.
    Thus, we assume that $n\geq 10$. It follows that $p_{1}+1=\alpha_{p}+4\leq m+4\leq n-1$. If $p_{1}+1=n-1$, then $\alpha_{p}=m$ and $F^{c}=\{m,m+3,m+4\}$. By Proposition 3.15, it follows that $F$ is not a spanning facet, a contradiction. Hence $p_{1}+1<n-1$ and thus, $p_{1}+1\in F$. We have $\alpha_{p}<_{\mathcal{O}}p_{1}+1$ and $p_{1}<p_{1}+1$. Let $F^{\prime c}:=(F^{c}\setminus\{p_{2}\})\sqcup\{p_{1}+1\}$. If $F^{\prime}\notin M(\Delta_{3}(C_{n}^{2}))$, then Proposition 3.19 implies that $F$ is not a spanning facet, which is a contradiction. So $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Then $F^{\prime c}=\{\alpha_{p}\}\sqcup\{\alpha_{p}+3,\alpha_{p}+4\}$. This means that $F^{\prime}$ satisfies $(\mathcal{D}_{4})$, and thus $F^{\prime}\in\mathcal{D}$. Hence, Proposition 3.19 implies that $F$ is not a spanning facet, again a contradiction. So $p_{1}\neq\alpha_{p}+3\text{ (mod \ $n$})$.
    Therefore, $p_{1}\in V(C_{n}^{2})\setminus(\{\alpha_{p}-4,\alpha_{p}-3,\alpha_{p}-2,\alpha_{p}-1,\alpha_{p},\alpha_{p}+1,\alpha_{p}+2,\alpha_{p}+3\ (\text{mod \ $n$})\}\sqcup\{y\})$.
@@
    We show that $F$ satisfies $(\mathcal{S}_{3})$. For this, we show that $p_{1}\in V(C_{n}^{2})\setminus(\{m-3,m-2,m-1,m,m+1,m+2,m+3,m+4\ (\text{mod $n$})\}\sqcup\{y\})$ (as $\min\{2m-\alpha_{p},\alpha_{p}-4\}=\alpha_{p}-4=m-3$). We have $F^{c}=\{m+1\}\sqcup\{p_{1},p_{2}\}$, where $p_{2}=n-1$.
    Suppose $p_{1}=y$. Since $p_{1}\neq p_{2}=n-1$, we get $y\in\{0,1\}$. If $y=0$, then $m+1=n-4$. Hence $p_{2}=n-1=m+4$, which implies that $F^{c}=\{m+1,(m+1)+3\text{ (mod \ $n$}),(m+1)+4\text{ (mod \ $n$})\}$. Thus, $F$ is not a spanning facet by Proposition 3.15. This is a contradiction. So $y=1$, and then $m+1>n-4$. This implies that $C_{n}^{2}[F^{c}]$ is connected, again a contradiction. Hence $p_{1}\neq y$.
@@
      Here, $p_{1}+1=m-2\in F$. We have $\alpha_{p}=m+1<_{\mathcal{O}}m-2=p_{1}+1$ and $p_{1}<p_{1}+1$. Observe that if $F^{\prime c}=\{m+1\}\sqcup\{m-3,m-2\}$, then $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$ such that $F^{\prime}$ satisfies $(\mathcal{D}_{3})$ and thus, $F^{\prime}\in\mathcal{D}$. Therefore, Proposition 3.19 implies that $F$ is not a spanning facet, a contradiction. Hence $p_{1}\neq m-3$.
@@
      We have $p_{1}+4=m+2\in F$. Also, $\alpha_{p}=m+1<_{\mathcal{O}}m+2=p_{1}+4$ and $p_{1}<p_{1}+4$. If $F^{\prime c}=\{m+1\}\sqcup\{m-2,m+2\}$, then $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$ such that $F^{\prime}$ satisfies $(\mathcal{D}_{1})$. Thus $F^{\prime}\in\mathcal{D}$. Therefore, by Proposition 3.19, $F$ is not a spanning facet. This is a contradiction. Hence $p_{1}\neq m-2$.
@@
      Observe that if $n=9$, then $n-1=m+3$; if $n\in\{10,11\}$, then $n-1=m+4$; and if $n\in\{12,13\}$, then $n-1=m+5$. We have $F^{c}=\{m+1\}\sqcup\{p_{1},n-1\}$. Since $C_{n}^{2}[F^{c}]$ is disconnected and $p_{1}<n-1$, it follows that if $p_{1}=m+2$, then $n\geq 12$; and if $p_{1}=m+3$, then $n\geq 14$. Therefore, in each case, we have $p_{1}+2<n-1$. This implies that $p_{1}+2\in F$. We have $\alpha_{p}<_{\mathcal{O}}p_{1}+2$, $p_{1}<p_{1}+2$ and $C_{n}^{2}[(F^{c}\setminus\{n-1\})\sqcup\{p_{1}+2\}]=C_{n}^{2}[\{\alpha_{p},p_{1},p_{1}+2\}]$ is connected. Using Proposition 3.19, we get a contradiction to the fact that $F$ is a spanning facet. Hence $p_{1}\notin\{m+2,m+3\}$.